### v1

# Pure C*-algebras

Abstract.
We show that every $C^{*}$-algebra that is $(m,n)$-pure in the sense of Winter is already pure, that is, its Cuntz semigroup is almost unperforated and almost divisible. More generally, we show that even weaker comparison and divisibility properties automatically lead to pureness.
As an application to the non-simple Toms-Winter conjecture, we show that every $C^{*}$-algebra with the Global Glimm Property and finite nuclear dimension is pure.
Key words and phrases:
$C^{*}$-algebras, Cuntz semigroups, comparison, divisibility, Global Glimm Property
2010 Mathematics Subject Classification:
Primary 46L05; Secondary 19K14, 46L80, 46L85.
RA and FP were partially supported by Spanish Research State Agency (grant No.  PID2020-113047GB-I00/AEI/10.13039/501100011033), by the Comissionat per Universitats i Recerca de la Generalitat de Catalunya (grant No. 2021-SGR-01015) and by the Spanish State Research Agency through the Severo Ochoa and María de Maeztu Program for Centers and Units of Excellence in R&D (CEX2020-001084-M). HT was partially supported by the Knut and Alice Wallenberg Foundation (KAW 2021.0140).
1. Introduction
The groundbreaking classification of amenable factors by Connes and Haagerup [Con76, Haa87] in the 1970s and 80s sparked the Elliott program to classify simple, nuclear $C^{*}$-algebras. A first breakthrough in this direction was obtained with the classification in the purely infinite case by Kirchberg and Phillips [Kir95, Phi00] in the 1990s. In the stably finite case, the game-changing examples of Toms [Tom08] showed that additional regularity properties have to be assumed for $K$-theoretic classification, and building on decades of research by numerous mathematicians, in a series of spectacular results [Win14, EGLN15, TWW17, GLN20], the classification of simple, nuclear $C^{*}$-algebras that are $\mathcal{Z}$-stable and satisfy the universal coefficient theorem (UCT) was accomplished in the 2010s. (See also the recent approach [CGS+23] and the survey [Whi23].)
While the UCT holds in all cases of interest (and possibly it is automatic for all nuclear $C^{*}$-algebras), it is known that $\mathcal{Z}$-stability can fail for simple, nuclear $C^{*}$-algebras, and it becomes highly relevant to determine when precisely this regularity property holds. To investigate this problem, Winter [Win12] introduced the notion of (m,n)-pureness for simple $C^{*}$-algebras, with $m$ and $n$ quantifying certain comparison and divisibility properties among operators, and he showed that every simple $C^{*}$-algebra that has locally finite nuclear dimension (a strong form of nuclearity) and that is $(m,n)$-pure for some $m,n$ is $\mathcal{Z}$-stable. Building on work of Rørdam [Rør04], Winter also showed that every simple, $\mathcal{Z}$-stable $C^{*}$-algebra is $(0,0)$-pure (abbreviated to just pure), which means that it satisfies the strongest of the comparison and divisibility properties.
The concept of $(m,n)$-pureness does not depend on the ideal structure of the algebra, hence may be stated for non-simple $C^{*}$-algebras; see Definition 5.1. However, as we show in this paper, all versions of pureness coincide. Our first main result is (see Theorem 5.8):
Theorem A.
If a $C^{*}$-algebra is $(m,n)$-pure for some $m,n\in{\mathbb{N}}$, then it is pure.
This should be contrasted with the well-known examples of simple, amenable $C^{*}$-algebras that do not have 00-comparison ([Vil98]), or that do not have 00-almost divisibility ([DHTW09]). Thus Theorem A can be interpreted as saying that if a $C^{*}$-algebra is not pure, then it necessarily has poor comparison or poor divisibility properties.
Previously, Theorem A has been obtained for separable, simple, unital $C^{*}$-algebras with locally finite nuclear dimension [Win12], and recently for arbitrary simple $C^{*}$-algebras [APRT24, Theorem 10.5]. Our proof of the general non-simple, non-nuclear case is based on the methods developed in [APRT24] in combination with new techniques about what we term functional divisibility (see Section 4 for details). In turn, these are partly inspired by results in [RT17].
Thus a key ingredient in our approach is the Cuntz semigroup and its technical features (see Section 2 for the precise definition). This is an invariant constructed by Cuntz in [Cun78], which is modelled after the Murray-von Neumann semigroup of projections but using positive elements instead. The version we use is what is commonly called the complete Cuntz semigroup, considered in [CEI08], and technically more suitable as it behaves well with respect to, for example, inductive limits. The power of the Cuntz semigroup was demonstrated in the context of the classification program: A. S. Toms produced in [Tom08] two $C^{*}$-algebras that were indistinguishable by the Elliott invariant and a number of reasonable extensions of it, but could be told apart using their Cuntz semigroup. The salient distinguishing feature was strict comparison of positive elements. Another, more recent use of this semigroup is related to the structure of $C^{*}$-algebras of stable rank one, carried out in [APRT22], where powerful Cuntz semigroup techniques yielded solutions to three structural open problems.
The Toms-Winter conjecture
This prominent conjecture [Win18] aims at establishing a robust class of $C^{*}$-algebras covered by the classification theorem referred to above. It predicts that for a separable, simple, unital, non-elementary, nuclear $C^{*}$-algebra, the following properties are equivalent:
  1. (1)
    $A$ has finite nuclear dimension.
  2. (2)
    $A$ is $\mathcal{Z}$-stable.
  3. (3)
    $A$ has strict comparison of positive elements.
The nuclear dimension was introduced in [WZ10] as a noncommutative generalization of covering dimension. A $C^{*}$-algebra $A$ is said to be $\mathcal{Z}$-stable if it tensorially absorbs the Jiang-Su algebra $\mathcal{Z}$, that is, if $A\cong\mathcal{Z}\otimes A$. A $C^{*}$-algebra $A$ has strict comparison of positive elements if it has 00-comparison; see Definitions 3.1 and 5.1 for details.
In two groundbreaking papers [Win12, CET+21], it was shown that conditions (1) and (2) of the conjecture are equivalent. This was later generalized to stably projectionless, simple $C^{*}$-algebras in [Tik14, CE20], thereby showing that a separable, simple, non-elementary $C^{*}$-algebra has finite nuclear dimension if and only if it is nuclear and $\mathcal{Z}$-stable.
Further, it is known that (2) implies (3) [Rør04]. The implication (3$\Rightarrow$(2) has been verified under certain additional assumptions [Sat12, KR14, TWW15, Thi20], but it remains open in general.
The Toms-Winter conjecture is also closely related to pureness, which we view as a property lying between (2) and (3). Indeed, as mentioned above, $\mathcal{Z}$-stability implies pureness, which in turn implies strict comparison of positive elements; see Proposition 5.2. Further, among separable, simple, unital, non-elementary $C^{*}$-algebras, pureness implies $\mathcal{Z}$-stability under the additional assumption of locally finite nuclear dimension [Win12], and strict comparison of positive elements implies pureness under the additional assumption of a positive solution to the Rank Problem. Such a positive solution has been obtained for $C^{*}$-algebras of stable rank one [Thi20, APRT22].
The non-simple Toms-Winter conjecture
The regularity properties of the Toms-Winter conjecture are also relevant for non-simple $C^{*}$-algebras, and an investigation of their interplay was initiated in [RT17]. While the assumption of unitality in the original formulation of the conjecture could be dispensed with, the condition to be non-elementary is crucial. Indeed, elementary $C^{*}$-algebras have nuclear dimension zero, but they are neither $\mathcal{Z}$-stable nor pure.
There are two natural generalizations of non-elementariness to the non-simple setting: Nowhere scatteredness, and the Global Glimm property. Here, a $C^{*}$-algebra is nowhere scattered if none of its quotients has a nonzero, elementary ideal, or equivalently if every hereditary sub-$C^{*}$-algebra is generated by nilpotent elements [TV24]. A $C^{*}$-algebra has the Global Glimm Property if every hereditary sub-$C^{*}$-algebra contains an approximately full, nilpotent element [TV23]. Clearly, every $C^{*}$-algebra with the Global Glimm Property is nowhere scattered, and the Global Glimm Problem asks whether the converse holds. We ask:
Question B.
Let $A$ be a separable, nowhere scattered, nuclear $C^{*}$-algebra. Are the following properties equivalent?
  1. (1)
    $A$ has finite nuclear dimension.
  2. (2)
    $A$ is $\mathcal{Z}$-stable.
  3. (3a)
    $A$ is pure.
  4. (3b)
    $A$ has strict comparison of positive elements.
The implications (2$\Rightarrow$(3a)$\Rightarrow$(3b) hold in general; see Proposition 5.2. We also note that pureness implies the Global Glimm Property; see [TV23, Theorem 3.6]. Therefore, Question B contains in particular the question of whether the Global Glimm Problem has a positive solution for $C^{*}$-algebras with finite nuclear dimension. As for simple $C^{*}$-algebras, the question whether strict comparison of positive elements implies pureness is closely related to the Rank Problem. This was solved for $C^{*}$-algebras of real rank zero or stable rank one ([ER06] and [APRT22]), which shows that (3b) implies (3a) in these cases.
In [RT17, Theorems 7.10, 7.15] it is shown that (3a) implies (2) under the additional assumption that the $C^{*}$-algebra $A$ has locally finite nuclear dimension, that no quotient of $A$ has a simple, purely infinite ideal, and that the primitive ideal space of $A$ has a basis of compact-open subsets or is Hausdorff.
The following result (which is Theorem 6.4) provides a further partial answer to Question B.
Theorem C.
Every $C^{*}$-algebra with the Global Glimm Property and finite nuclear dimension is pure.
Acknowledgements
The authors thank Leonel Robert for valuable feedback on an earlier version of this paper.
2. Preliminaries
The techniques used in the paper mainly concern the structure of the Cuntz semigroup $\operatorname{Cu}(A)$ of a $C^{*}$-algebra $A$, whose definition and properties we describe below.
Given positive elements $a,b$ in a $C^{*}$-algebra $A$, one says that $a$ is Cuntz subequivalent to $b$, denoted by $a\precsim b$, if there is a sequence $(r_{n})_{n}$ in $A$ such that $\lim_{n\to\infty}\|a-r_{n}br_{n}^{*}\|=0$. Further, $a$ and $b$ are Cuntz equivalent, denoted $a\sim b$, if $a\precsim b$ and $b\precsim a$. These relations were introduced by Cuntz in [Cun78].
[…]
It is proved in [CEI08] that $\operatorname{Cu}(A)$ satisfies the following properties defined for a general positively ordered monoid $S$ (and usually referred to as axioms):
  1. (O1)
    If $(x_{n})_{n}$ is an increasing sequence in $S$, then $\sup_{n}x_{n}$ exists.
  2. (O2)
    For any $x\in S$ there exists a sequence $(x_{n})_{n}$ such that $x_{n}\ll x_{n+1}$ for all $n$ and $x=\sup_{n}x_{n}$. (We say that $(x_{n})_{n}$ is a $\ll$-increasing sequence.)
  3. (O3)
    If $x_{1}\ll x_{2}$ and $y_{1}\ll y_{2}$, then $x_{1}+y_{1}\ll x_{2}+y_{2}$.
  4. (O4)
    If $(x_{n})_{n}$ and $(y_{n})_{n}$ are increasing sequences in $S$, then $\sup_{n}(x_{n}+y_{n})=\sup_{n}x_{n}+\sup_{n}y_{n}$.
The relation $\ll$ in these axioms is defined as follows: $x\ll y$ if for every increasing sequence $(y_{n})_{n}$ satisfying $y\leq\sup_{n}y_{n}$ there exists $n_{0}\in{\mathbb{N}}$ such that $x\leq y_{n_{0}}$. The relation $\ll$ is called the way-below relation, or compact containment relation, and one says that ‘$x$ is way-below $y$’ if $x\ll y$. An element $u\in S$ such that $u\ll u$ is termed compact.
A positively ordered monoid satisfying axioms (O1)-(O4) is called a $\mathrm{Cu}$-semigroup and belongs to a category of semigroups that has thoroughly been studied (see, for example, [APT18], [APT20a], [APT20b] [APT20c], [TV22], [APRT22], among others; see also the recent survey [GP23]). Yet, when dealing with semigroups obtained as Cuntz semigroups of $C^{*}$-algebras, or further restricting to particular families of $C^{*}$-algebras, additional axioms and properties are satisfied.
The first axiom that is added after (O4) is the following:
  1. (O5)
    For all $x^{\prime},x,y$ with $x^{\prime}\ll x\leq y$ there exists $z$ such that $x^{\prime}+z\leq y\leq x+z$. Moreover, if $x+w\leq y$ for some $w$, and $w^{\prime}\ll w$, then $z$ may be chosen such that $w^{\prime}\ll z$.
This axiom is usually referred to as the axiom of almost algebraic order. The first part of the statement above was proved in [RW10, Lemma 7.1], and the second part can be found in [APT18, Proposition 4.6].
Since we will often use the following consequence of (O5), we add a proof here for completeness.
Lemma 2.1 ([TV23, Lemma 2.2]).
Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5), let $k\in{\mathbb{N}}$, and let $z^{\prime},z,x\in S$ satisfy $z^{\prime}\ll z$ and $(k+1)z\leq x$. Then there exists $d\in S$ such that
$kz^{\prime}+d\leq x\leq kz+d,\,\,\,\text{ and }\,\,\,x\leq(k+1)d.$
Proof.
Choose $z^{\prime\prime}\in S$ such that $z^{\prime}\ll z^{\prime\prime}\ll z$. Applying (O5) for
$kz^{\prime}\ll kz^{\prime\prime},\quad z^{\prime\prime}\ll z,\,\,\,\text{ and }\,\,\,kz^{\prime\prime}+z\leq x,$
we obtain $d\in S$ satisfying
$kz^{\prime}+d\leq x\leq kz^{\prime\prime}+d,\,\,\,\text{ and }\,\,\,z^{\prime\prime}\ll d.$
Then $x\leq(k+1)d$, which shows that $d$ satisfies the desired conditions. ∎
Two more axioms are used very often in the theory of Cu-semigroups. The first one, (O6), is a weaker form of Riesz decomposition and the second one, (O7), is a weak form of interpolation. That the Cuntz semigroup of any $C^{*}$-algebra satisfies both (O6) and (O7) was proved in [Rob13, Proposition 5.1.1] and [APRT21, Proposition 2.2], respectively. We will not need the precise formulation of these axioms here, but rather a very useful consequence of them that allows to take infima with idempotent elements. We make this precise below.
Recall that a $\mathrm{Cu}$-semigroup $S$ is said to be countably based if all elements can be written as suprema of elements from a fixed countable subset. It is known that $\operatorname{Cu}(A)$ is countably based whenever $A$ is separable (for a proof, see [APS11, Lemma 1.3]).
Theorem 2.2 (cf. [APRT21, Theorems 2.4, 2.5]).
Let $S$ be a countably based $\mathrm{Cu}$-semigroup satisfying (O5)-(O7). Then each $x\in S$ and each idempotent element $w\in S$ have an infimum $x\wedge w$ in $S$. Further, the map $S\to S$ given by $x\mapsto x\wedge w$ is a monoid homomorphism preserving the order and the suprema of increasing sequences.
Let $S$ be a $\mathrm{Cu}$-semigroup. A map $\lambda\colon S\to[0,\infty]$ is called a functional if $\lambda$ preserves addition, the zero element, order and suprema of increasing sequences. We denote by $F(S)$ the set of functionals on $S$. This is a cone when endowed with the operations of pointwise addition and pointwise scalar multiplication by positive real numbers. When equipped with the appropriate topology, $F(S)$ becomes a compact Hausdorff cone (see [ERS11, Theorems 4.4, 4.8], [Rob13, Section 2.2]; see also [Kei17, Theorem 3.17]).
Given $x\in S$, we obtain a map $\widehat{x}\colon F(S)\to[0,\infty]$ defined by $\widehat{x}(\lambda)=\lambda(x)$, for $\lambda\in F(S)$. $\widehat{x}$ is linear and lower semicontiuous, and if $S$ satisfies (O5) and (O6), then $\textrm{Lsc}(F(S))$, the set of all such maps, is an algebraically ordered, lattice ordered, compact cone.
Two main questions arise in this setting. To determine the range of the so-called rank map $x\mapsto\widehat{x}$ (referred to as the Rank Problem); and determine to what extent the order in $S$ can be inferred by the order in $\textrm{Lsc}(F(S))$. Note that $x\mapsto\widehat{x}$ is additive and preserves sequential suprema; it does not preserve the relation of compact containment, and is certainly not an order embedding. However, the result below will be useful.
Given elements $x$ and $y$ in a partially ordered semigroup, one writes $x<_{s}y$ if there exists $k\in{\mathbb{N}}$ such that $(k+1)x\leq ky$. Item (ii) in the following lemma is a consequence of [Rob13, Proposition 2.2.2] or [APT18, Theorem 5.2.19].
Proposition 2.3 ([Rob13, Lemma 2.2.5], [APRT24, Lemma 5.3]).
Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5), and let $x,y\in S$. Then:
  1. (i)
    If $x\ll y$, and $s,t\in(0,\infty]$ satisfy $s<t$, we have $s\widehat{x}\ll t\widehat{y}$ in $\operatorname{Lsc}(F(S))$.
  2. (ii)
    If $\widehat{x}\leq\gamma\widehat{y}$ for some $\gamma\in(0,1)$, then $x^{\prime}<_{s}y$ for all $x^{\prime}\ll x$.
3. Variants of comparison
We first recall the definition of $m$-comparison as considered by Toms and Winter in [TW09, Lemma 6.1], and later formalized by Ortega, Perera and Rørdam in [OPR12, Definition 2.8].
Definition 3.1.
Let $m\in{\mathbb{N}}$. A $\mathrm{Cu}$-semigroup $S$ has $m$-comparison if for all $x,y_{0},\ldots,y_{m}\in S$ satisfying $x<_{s}y_{j}$ for $j=0,\ldots,m$, we have $x\leq y_{0}+\ldots+y_{m}$. The lowest possible value for $m$-comparison is attained when $m=0$, which is commonly referred to as saying that $S$ is almost unperforated.
We say that a $C^{*}$-algebra has $m$-comparison if its Cuntz semigroup does.
Remarks 3.2.
(1) L. Robert showed in [Rob11, Theorem 1.3] that, if $A$ is a $C^{*}$-algebra with nuclear dimension at most $m$, then $A$ has $m$-comparison.
(2) Let $m\in{\mathbb{N}}$ and let $S$ be a $\mathrm{Cu}$-semigroup that has $m$-comparison. Then, $x<_{s}y$ implies $x\leq(m+1)y$, for all $x,y\in S$.
Given $N\in{\mathbb{N}}$ and elements $x$ and $y$ in a partially ordered semigroup, we write $x\leq_{N}y$ if $nx\leq ny$ for all $n\in{\mathbb{N}}$ with $n\geq N$; see [APRT24, Notation 5.2] and also [APT18, Definition 5.6.1]. To recall the definition of locally bounded comparison amplitude, we need the notion of a scale in a $\mathrm{Cu}$-semigroup.
3.3Scales.
Let $S$ be a $\mathrm{Cu}$-semigroup. A scale in $S$ is a subset $\Sigma\subseteq S$ such that
  1. (i)
    is downward hereditary: if $x\leq y$ for some $x\in S$ and $y\in\Sigma$, then $x\in\Sigma$,
  2. (ii)
    is closed under suprema of increasing sequences, and
  3. (iii)
    $\Sigma$ generates $S$ as an ideal, that is, for $x^{\prime},x\in S$ with $x^{\prime}\ll x$, there exist $y_{1},\ldots,y_{n}\in\Sigma$ such that $x^{\prime}\leq y_{1}+\ldots+y_{n}$.
(See [APT20c, Definition 4.1]). We say that the pair $(S,\Sigma)$ is a scaled $\mathrm{Cu}$-semigroup.
Given a scale $\Sigma$ and $d\in{\mathbb{N}}$, the $d$-fold amplification of $\Sigma$ is defined as $\Sigma^{(0)}:=\{0\}$, and
[…]
for $d\geq 1$.
If $A$ is a $C^{*}$-algebra, then
$\Sigma_{A}:=\big{\{}x\in\operatorname{Cu}(A):x\leq[a]\text{ for some $a\in A_{+}$}\big{\}}$
is a scale in $\operatorname{Cu}(A)$, and the pair $(\operatorname{Cu}(A),\Sigma_{A})$ is called the scaled Cuntz semigroup of $A$; see [APT20c, Paragraph 4.2] and [TV23, Lemma 3.3], and also [APRT24, Paragraph 4.5]. By [APRT24, Proposition 7.4], for $d\geq 1$ we have
$\Sigma_{A}^{(d)}=\big{\{}x\in\operatorname{Cu}(A):x\leq[a]\text{ for some $a\in M_{d}(A)_{+}$}\big{\}}.$
We recall below the definition of locally bounded comparison amplitude, from [APRT24]. Useful reformulations of this condition will be given in Lemma 3.5.
Definition 3.4 ([APRT24, Definition 6.1]).
A scaled $\mathrm{Cu}$-semigroup $(S,\Sigma)$ has locally bounded comparison amplitude, in short (LBCA), if for every $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$ there exists $N=N(\gamma,d)$ such that the following holds:
$\widehat{x}\leq\gamma\widehat{y}\quad\textrm{ implies }\quad x\leq_{N}y,\quad\text{ for all }x,y\in\Sigma^{(d)}.$
A $C^{*}$-algebra will be said to have locally bounded comparison amplitude if $(\operatorname{Cu}(A),\Sigma_{A})$ has (LBCA).
Lemma 3.5.
Let $(S,\Sigma)$ be a scaled $\mathrm{Cu}$-semigroup. The following are equivalent:
  1. (i)
    $(S,\Sigma)$ has locally bounded comparison amplitude.
  2. (ii)
    For every $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$ there exists $N=N(\gamma,d)$ such that
    $\widehat{x}\leq\gamma\widehat{y^{\prime}},\quad y^{\prime}\ll y\quad\text{ implies }\quad Nx\leq Ny,\qquad\text{for all }x,y^{\prime},y\in\Sigma^{(d)}.$
  3. (iii)
    For every $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$ there exists $M=M(\gamma,d)$ such that:
    $\widehat{x}\leq\gamma\widehat{y}\quad\text{ implies }\quad(M+1)x\leq My,\qquad\text{for all }x,y\in\Sigma^{(d)}.$
Proof.
It is clear that (i) implies (ii).
Let us show that (ii) implies (iii). Fix $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$. Choose $\gamma^{\prime}$ and $n\in{\mathbb{N}}$ such that
$\gamma<\gamma^{\prime},\quad\frac{n+1}{n}\gamma^{\prime}<1.$
Let $N=N(\tfrac{n+1}{n}\gamma^{\prime},d(n+1))$ as provided by (ii). We claim that $M=Nn$ has the desired properties to verify (iii) (for $\gamma$ and $d$).
To this end, let $x,y\in\Sigma^{(d)}$ satisfy $\widehat{x}\leq\gamma\widehat{y}$. Let $x^{\prime}\in S$ be such that $x^{\prime}\ll x$. By Proposition 2.3 (i) applied to $x^{\prime}\ll x$ and $1<\frac{\gamma^{\prime}}{\gamma}$, we have that $\widehat{x^{\prime}}\ll\gamma^{\prime}\widehat{y}$. Hence, there exists $y^{\prime}\ll y$ such that $\widehat{x^{\prime}}\leq\gamma^{\prime}\widehat{y^{\prime}}$. Then
$(n+1)\widehat{x^{\prime}}\leq(n+1)\gamma^{\prime}\widehat{y^{\prime}}=\left(\frac{n+1}{n}\gamma^{\prime}\right)n\widehat{y^{\prime}}.$
Since $(n+1)x^{\prime},ny^{\prime}$ and $ny$ belong to $\Sigma^{((n+1)d)}$, it follows from the choice of $N$ that $N(n+1)x^{\prime}\leq Nny$. Passing to the supremum over all $x^{\prime}\ll x$ we get that $N(n+1)x\leq Nny$, and thus $(M+1)x\leq My$.
Finally, note that if elements $x$ and $y$ in a partially ordered semigroup satisfy $(M+1)x\leq My$ for some $M\in{\mathbb{N}}$, then there exists $N\in{\mathbb{N}}$ such that $(n+1)x\leq ny$ for all $n\geq N$, and in particular $x\leq_{N}y$; see [APT18, Proposition 5.2.13]. This shows that (iii) implies (i). ∎
The following lemma is implied by [APRT24, Proposition 10.3].
Lemma 3.6.
Any $C^{*}$-algebra with $m$-comparison has locally bounded comparison amplitude.
Proof.
In order to apply [APRT24, Proposition 10.3] it is required that the Cuntz semigroup of a $C^{*}$-algebra satisfies (O5), (O6), and the so-called Edward’s condition. The first two are always satisfied, as mentioned in the preliminaries. As for the third condition, this was also shown to hold in general in [APRT21, Theorem 5.3]. ∎
Theorem 3.7.
If a $C^{*}$-algebra has finite nuclear dimension, then its Cuntz semigroup has locally bounded comparison amplitude.
Proof.
Let $A$ be a $C^{*}$-algebra with nuclear dimension at most $m$ for some $m\in{\mathbb{N}}$. As observed in Remarks 3.2, $\operatorname{Cu}(A)$ has $m$-comparison, hence the result follows from Lemma 3.6. ∎
4. Variants of divisibility
We first recall the definition of $n$-almost divisibility for $\mathrm{Cu}$-semigroups, which appeared in [RT17, Section 2.3]. Other related notions of divisibility will be also useful in our analysis. These have sprung into prominence due to their connection with important structural properties of $C^{*}$-algebras.
Definition 4.1.
Let $k\in{\mathbb{N}}$, and $n\in{\mathbb{N}}\cup\{\infty\}$. We say that an element $x$ in a $\mathrm{Cu}$-semigroup $S$ is $(k,n)$-divisible if, for each $x^{\prime}\in S$ with $x^{\prime}\ll x$, there exists $y\in S$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\leq(k+1)(n+1)y.$
If all elements are $(k,n)$-divisible, when $n=\infty$, then we say that $S$ is $(k,\omega)$-divisible.
If all elements are $(k,n)$-divisible for all $k$, we then say that $S$ is $n$-almost divisible. In case $S$ is 00-almost divisible, then we just say that $S$ is almost divisible.
A $C^{*}$-algebra is $n$-almost divisible if its Cuntz semigroup is.
Remarks 4.2.
(1) It was proved in [TV24, Theorem 9.1] that, if $A$ is a nowhere scattered $C^{*}$-algebra of real rank zero, then $A$ is almost divisible. In fact, its Cuntz semigroup $\operatorname{Cu}(A)$ is weakly divisible: for any $x\in\operatorname{Cu}(A)$, there are $y,z\in\operatorname{Cu}(A)$ such that $x=2y+3z$; see also [APGPSM10, Section 6]. In the simple case, and for the Murray-von Neumann semigroup, this was partly obtained in [PR04, Proposition 5.3].
(2) If a $\mathrm{Cu}$-semigroup $S$ is almost unperforated, then it is almost divisible precisely when, given $x\in S$ and $k$, there is $y\in S$ such that $ky\leq x\leq(k+1)y$. This was proved in [APT18, Proposition 7.3.7].
(3) The concept of $(k,\omega)$-divisibility is connected to the Global Glimm Property which, in the context of this paper, will be analysed with more detail in Section 6. It was proved in [TV23, Theorem 3.6] that a $C^{*}$-algebra has the Global Glimm Property precisely when its Cuntz semigroup is $(2,\omega)$-divisible.
Inspired by the results in [RT17, Section 6], we introduce the following more general divisibility conditions in terms of comparison by functionals.
Definition 4.3.
We say that an element $x$ in a $\mathrm{Cu}$-semigroup $S$ is
  • (i)
    functionally almost divisible, or that $x$ has (FAD), if for every $x^{\prime}\in S$ satisfying $x^{\prime}\ll x$, every $k\in{\mathbb{N}}_{>0}$, and every $\varepsilon>0$, there exists $y\in S$ such that
    $ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq k\widehat{y}+\varepsilon\widehat{x}.$
  • (ii)
    lower functionally almost divisible, or that $x$ has (LFAD), if for every $x^{\prime}\in S$ satisfying $x^{\prime}\ll x$, and every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ such that
    $ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq\left(k+1\right)\widehat{y}.$
  • (iii)
    upper functionally almost divisible, or that $x$ has (UFAD), if for every $x^{\prime}\in S$ satisfying $x^{\prime}\ll x$, and every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ and $\nu>0$ such that
    $y\leq x,\quad(k+\nu)\widehat{y}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x^{\prime}\leq(k+1)y.$
A $\mathrm{Cu}$-semigroup $S$ is said to be (upper, lower) functionally almost divisible if every element in $S$ is (upper, lower) functionally almost divisible.
We begin our study of functional divisibility by providing a number of technical results relating each of the notions above. These results will be applied to $C^{*}$-algebras with $n$-almost divisibility at the end of the section, where we will show that any $n$-almost divisible $C^{*}$-algebra satisfies all three divisibility notions; see Theorem 4.7.
Lemma 4.4.
Let $S$ be a $\mathrm{Cu}$-semigroup.
  1. (i)
    $S$ is functionally almost divisible if and only if for every $x^{\prime},x\in S$ satisfying $x^{\prime}\ll x$, and every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ such that
    $ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq k\widehat{y}+\tfrac{1}{k}\widehat{x}.$
  2. (ii)
    $S$ is lower functionally almost divisible if and only if for every $x^{\prime},x\in S$ satisfying $x^{\prime}\ll x$, and every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ such that
    $ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq(k+\tfrac{1}{k})\widehat{y}.$
Proof.
We will only prove (i) since the proof of (ii) is similar. Moreover the forward implications are clear. To prove the backwards implication, let $x^{\prime},x\in S$ satisfy $x^{\prime}\ll x$, let $k\in{\mathbb{N}}_{>0}$, and let $\varepsilon>0$. Choose $n\in{\mathbb{N}}_{>0}$ such that $\tfrac{1}{kn}\leq\varepsilon$. Applying the assumption for $x^{\prime},x$ and $kn$, we obtain $z\in S$ such that
$knz\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq kn\widehat{z}+\frac{1}{kn}\widehat{x}.$
Set $y:=nz$. Then
$ky=knz\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq kn\widehat{z}+\frac{1}{kn}\widehat{x}=k\widehat{y}+\frac{1}{kn}\widehat{x}\leq k\widehat{y}+\varepsilon\widehat{x},$
as desired. ∎
Lemma 4.5.
Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5) and (O7). Assume that $S$ is $(2,\omega)$-divisible and is functionally almost divisible. Then $S$ is lower functionally almost divisible.
Proof.
Let $k\in{\mathbb{N}}$, and let $x^{\prime},x\in S$ be such that $x^{\prime}\ll x$. It suffices to prove that there exists $y\in S$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq\left(k+\frac{1}{k}\right)\widehat{y}.$
For $k=1$ use $y:=x$. Thus, we may from now on assume that $k\geq 2$.
Using the methods from [TV21, Section 5], we find a countably based, $(2,\omega)$-divisible $\mathrm{Cu}$-semigroup $H$ that contains $x^{\prime},x$ and such that $H$ satisfies (O5), (O7), and (FAD).
Choose $x^{\prime\prime}\in H$ and $n\in{\mathbb{N}}_{>0}$ such that
$x^{\prime}\ll x^{\prime\prime}\ll x,\quad k+1\leq n,\,\,\,\text{ and }\,\,\,\frac{k(k+1)+n}{n}\leq 1+\frac{1}{k^{2}}.$
Since $H$ is $(2,\omega)$-divisible, it is also $(n,\omega)$-divisible by the arguments in [TV23, Lemma 3.4]. Applied to $x^{\prime\prime}\ll x$ and $n$, we obtain $c\in H$ such that
$nc\leq x,\,\,\,\text{ and }\,\,\,x^{\prime\prime}\ll\infty c.$
Using that $x^{\prime\prime}\ll\infty c$, we can choose $c^{\prime}\in H$ such that
$x^{\prime\prime}\ll\infty c^{\prime},\,\,\,\text{ and }\,\,\,c^{\prime}\ll c.$
We have $(k+1)c\leq nc\leq x$. Using Lemma 2.1, we obtain $d\in H$ such that
$kc^{\prime}+d\leq x\leq kc+d,\,\,\,\text{ and }\,\,\,x\leq(k+1)d.$
Thus, one has
$nx^{\prime\prime}\leq nx\leq nkc+nd\leq kx+nd\leq(k(k+1)+n)d.$
Since $H$ is countably based and satisfies (O7), we may apply Theorem 2.2. Hence, for every $s\in S$ the infimum $s\wedge\infty c^{\prime}$ exists, and the map $s\mapsto s\wedge\infty c^{\prime}$ is additive; Set $e:=d\wedge\infty c^{\prime}$. Since $x^{\prime\prime}\ll\infty c^{\prime}$, we have
$nx^{\prime}\ll nx^{\prime\prime}=(nx^{\prime\prime})\wedge(\infty c^{\prime})$ $\leq(k(k+1)+n)d\wedge(\infty c^{\prime})$
$=\big{(}k(k+1)+n\big{)}\big{(}d\wedge(\infty c^{\prime})\big{)}=(k(k+1)+n)e.$
This allows us to find $e^{\prime}\in H$ satisfying
$nx^{\prime}\leq(k(k+1)+n)e^{\prime},\,\,\,\text{ and }\,\,\,e^{\prime}\ll e.$
Let us further choose $e^{\prime\prime}$ such that $e^{\prime}\ll e^{\prime\prime}\ll e$.
There exists $N\in{\mathbb{N}}$ such that $e^{\prime\prime}\leq Nc^{\prime}$. Choose $\varepsilon>0$ small enough such that
$\left(1+\frac{1}{k^{2}}\right)\varepsilon\leq\frac{1}{N}.$
Applying (FAD) to $e^{\prime}\ll e^{\prime\prime}$, $k$ and $\varepsilon$, we obtain $f\in H$ such that
$kf\leq e^{\prime\prime},\,\,\,\text{ and }\,\,\,\widehat{e^{\prime}}\leq\widehat{kf}+\varepsilon\widehat{e^{\prime\prime}}.$
Set $y:=c^{\prime}+f$. We have $kf\leq e\leq d$ and therefore
$ky=kc^{\prime}+kf\leq kc^{\prime}+d\leq x.$
To verify that $\widehat{x^{\prime}}\leq(k+1/k)\widehat{y}$, let $\lambda\in F(H)$. Then
$\begin{split}\lambda(x^{\prime})&\leq\frac{k(k+1)+n}{n}\lambda(e^{\prime})\leq\left(1+\frac{1}{k^{2}}\right)\lambda(e^{\prime})\\ &\leq\left(1+\frac{1}{k^{2}}\right)\left(k\lambda(f)+\varepsilon\lambda(e^{\prime\prime})\right)\leq\left(k+\frac{1}{k}\right)\lambda(f)+\frac{1}{N}\lambda(e^{\prime\prime})\\ &\leq\left(k+\frac{1}{k}\right)\lambda(f)+\lambda(c^{\prime})\leq\left(k+\frac{1}{k}\right)\lambda(y),\end{split}$
as desired. ∎
Lemma 4.6.
Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5), and let $x\in S$ be a compact element. If $x$ is functionally almost divisible, then for every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ such that
$y\leq x,\quad(k+\tfrac{1}{2})\widehat{y}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x\leq(k+1)y.$
In particular, $x$ is upper functionally almost divisible.
Proof.
Let $k\in{\mathbb{N}}$. Using (FAD) with $k+1$ and $\varepsilon=1/3(2k+1)(k+1)$, we obtain an element $z\in S$ such that
$(k+1)z\leq x,\,\,\,\text{ and }\,\,\,\widehat{x}\leq(k+1)\widehat{z}+\frac{1}{3(2k+1)(k+1)}\widehat{x}.$
Find $\delta>0$ such that
$\frac{k+1}{1-\delta}\leq(k+1)+\frac{1}{2(2k+1)},\,\,\,\text{ and }\,\,\,\frac{\varepsilon}{(1-\delta)}\leq\frac{1}{2(2k+1)(k+1)}.$
Now, using that $x\ll x$, it follows from Proposition 2.3 (i) that
$(1-\delta)\widehat{x}\ll\widehat{x}\leq(k+1)\widehat{z}+\varepsilon\widehat{x}.$
This allows us to choose $z^{\prime}\ll z$ such that
(1) $\widehat{x}\leq\frac{k+1}{(1-\delta)}\widehat{z^{\prime}}+\frac{\varepsilon}{(1-\delta)}\widehat{x}\leq\left((k+1)+\frac{1}{2(2k+1)}\right)\widehat{z^{\prime}}+\frac{1}{2(2k+1)(k+1)}\widehat{x}$
Using Lemma 2.1, we can now find $y\in S$ such that
$kz^{\prime}+y\leq x\leq(k+1)y.$
Take $\lambda\in F(S)$. If $\lambda(z^{\prime})=\infty$, we have $\lambda(y)=\lambda(x)=\infty$ and so, in particular, $(k+\tfrac{1}{2})\lambda(y)\leq\lambda(x)$.
Thus, assume that $\lambda(z^{\prime})<\infty$. Since $k\lambda(y)\leq k(\lambda(x)-k\lambda(z^{\prime}))$, using (1) at the first step, and $(k+1)z^{\prime}\leq x$ at the last step, one has:
$\begin{split}(k+\tfrac{1}{2})\lambda(y)&\leq(k+\tfrac{1}{2})\left(\left((k+1)+\tfrac{1}{2(2k+1)}\right)\lambda(z^{\prime})+\tfrac{1}{2(2k+1)(k+1)}\lambda(x)-k\lambda(z^{\prime})\right)\\ &=(k+\tfrac{1}{2})\left(\left(1+\tfrac{1}{2(2k+1)}\right)\lambda(z^{\prime})+\tfrac{1}{2(2k+1)(k+1)}\lambda(x)\right)\\ &\leq(k+\tfrac{1}{2})\left(\tfrac{1}{k+1}\left(1+\tfrac{1}{2(2k+1)}\right)\lambda(x)+\tfrac{1}{2(2k+1)(k+1)}\lambda(x)\right)=\lambda(x),\end{split}$
as desired.
The last assertion is clear. ∎
Theorem 4.7.
Let $A$ be a $C^{*}$-algebra. Assume that $A$ is $n$-almost divisible for some $n\in{\mathbb{N}}$. Then $\operatorname{Cu}(A)$ is (upper, lower) functionally almost divisible.
Proof.
Assume that $A$ is $n$-almost divisible. Without loss of generality, we may assume $A$ to be stable. Further, note that we only need to prove that $\operatorname{Cu}(A)$ has (FAD) and (UFAD). Indeed, any $n$-almost divisible $C^{*}$-algebra has a $(2,\omega)$-divisible Cuntz semigroup. Thus, it follows from Lemma 4.5 that $\operatorname{Cu}(A)$ has (LFAD) whenever it has (FAD).
Let $\mathcal{U}$ be a free ultrafilter on ${\mathbb{N}}$, and let $a\in A_{+}$. Set $C=\{a\}^{\prime}\cap A_{\mathcal{U}}$ and $I=\{a\}^{\perp}\cap A_{\mathcal{U}}$. Then, it follows from [RT17, Theorem 6.1] that the class of the unit $[1]$ in $\operatorname{Cu}(C/I)$ has (FAD). Therefore, $[1]$ has (UFAD) by Lemma 4.6.
Now choose $k\in{\mathbb{N}}_{>0}$. Applying (FAD) to $[1]\ll[1]$ with $\varepsilon=\tfrac{1}{k+1}$, we obtain an element $e\in C/I$ such that
$k[e]\leq[1],\,\,\,\text{ and }\,\,\,\widehat{[1]}\leq k\widehat{[e]}+\tfrac{1}{k+1}\widehat{[1]}.$
Setting $\delta=\tfrac{1}{k(k+1)}$, we have $(1+\delta)\widehat{k[1]}\leq k^{2}\widehat{[e]}+\widehat{[1]}$. Now we apply Proposition 2.3(ii) to find a positive integer $s$ such that $(s+1)k[1]\leq s(k^{2}[e]+[1])$.
Separately, we apply (UFAD) to $[1]\ll[1]$ and we obtain an element $\tilde{f}\in C/I$ such that
$[\tilde{f}]\leq[1],\quad(k+\nu)\widehat{[\tilde{f}]}\leq\widehat{[1]},\,\,\,\text{ and }\,\,\,[1]\leq(k+1)[\tilde{f}],$
for some $10<\nu<$ 1. Since $[1]\ll[1]$, we can find $f\in C/I$ with $[f]\ll[\tilde{f}]$ such that $[1]\leq(k+1)[f]$. Again, applying Proposition 2.3(ii), we find a positive integer $t$ such that $(t+1)k[f]\leq t[1]$.
Let $\bar{e},\bar{f}\in C$ be lifts of $e,f$ respectively. Then, there exists $z\in\operatorname{Cu}(I)$ such that
$k[\bar{e}]\leq[1]+z,\quad(s+1)k[1]\leq s(k^{2}[\bar{e}]+[1])+z,$
and
$[\bar{f}]\leq[1]+z,\quad(t+1)k[\bar{f}]\leq t[1]+z,\quad[1]\leq(k+1)[\bar{f}]+z$
in $\operatorname{Cu}(C)$.
Set $b=a\bar{e}$ and $c=a\bar{f}$ in $A_{\mathcal{U}}$. Using that every element in $C$ commutes with $a$ (in particular, those elements applying the Cuntz subequivalences above) and that every element in $I$ is orthogonal to $a$, one obtains
$k[b]\leq[a],\quad(s+1)k[a]\leq s(k^{2}[b]+[a]),$
and
$[c]\leq[a],\quad(t+1)k[c]\leq t[a],\quad[a]\leq(k+1)[c]$
in $\operatorname{Cu}(A_{\mathcal{U}})$.
This shows that $[a]\in\operatorname{Cu}(A_{\mathcal{U}})$ has (FAD) and (UFAD). Now, using the fact that functional comparison can be restated in terms of comparison in the semigroup ([Rob13, Proposition 2.2.2]), a standard model theoretic argument shows that $[a]\in\operatorname{Cu}(A)$ also has (FAD) and (UFAD); see [FHL+21, Proposition 8.1.6] and also [APRT24, Theorem 10.5]. ∎
5. Pure C*-algebras
In this section, we show that every $(m,n)$-pure $C^{*}$-algebra is pure; see Theorem 5.8.
Definition 5.1.
Given $m,n\in{\mathbb{N}}$, we say that a $\mathrm{Cu}$-semigroup $S$ is $(m,n)$-pure if $S$ has $m$-comparison (Definition 3.1) and is $n$-almost divisible (Definition 4.1). We say that $S$ is pure if it is $(0,0)$-pure.
A $\mathrm{Cu}$-semigroup $S$ has strict comparison if, whenever $x,y\in S$ with $x\leq\infty y$ satisfy that $\lambda(x)<\lambda(y)$ for any $\lambda\in F(S)$ normalized at $y$, one has $x\leq y$.
A $C^{*}$-algebra is $(m,n)$-pure if its Cuntz semigroup is. Similarly, a $C^{*}$-algebra is pure if it is $(0,0)$-pure. We also say that a $C^{*}$-algebra has strict comparison of positive elements (by $2$-quasitraces) provided its Cuntz semigroup has strict comparison.
It was proved in [ERS11, Proposition 6.2] that $S$ has strict comparison if and only if it is almost unperforated.
Proposition 5.2.
Every $\mathcal{Z}$-stable $C^{*}$-algebra is pure. Every pure $C^{*}$-algebra has strict comparison of positive elements.
Proof.
Let $A$ be a $\mathcal{Z}$-stable $C^{*}$-algebra. It was shown in [Rør04, Theorem 4.5] that $\operatorname{Cu}(A)$ is almost unperforated, and thus by the comment above $A$ has strict comparison of positive elements. The argument used to prove almost unperforation can be used to show almost divisibility. For a proof, see e.g. [APT11, Theorem 5.35]. ∎
Remark 5.3.
It is perhaps worth pointing out that the results in [Rør04] deal with the uncomplete Cuntz semigroup $\mathrm{W}(A)$. This makes no difference regarding almost unperforation, since $\operatorname{Cu}(A)$ is a completion of $\mathrm{W}(A)$ (see [APT18, Theorem 3.2.8]), hence $\mathrm{W}(A)$ is almost unperforated if and only if $\operatorname{Cu}(A)$ is.
One can ask whether, for $\mathrm{Cu}$-semigroups, $(0,0)$-pureness agrees with some sort of tensorial absorption. This was explored in [APT18], together with an extensive analysis of the tensor product in the category $\operatorname{Cu}$.
Theorem 5.4 ([APT18, Theorem 7.3.11]).
A $\mathrm{Cu}$-semigroup $S$ is pure if and only if $S\cong\operatorname{Cu}(\mathcal{Z})\otimes S$.
If the Cuntz semigroup of a $C^{*}$-algebra $A$ is isomorphic to that of $\mathcal{Z}\otimes A$, then $A$ is clearly pure. In the proof of [Tom11, Theorem 1.2], Toms shows that the converse holds for simple $C^{*}$-algebras. (See also [APP18, Section 7].) The following question is thus pertinent:
Question 5.5.
Let $A$ be a pure $C^{*}$-algebra. Is $\operatorname{Cu}(A)\cong\operatorname{Cu}(\mathcal{Z}\otimes A)$?
We start with an important technical result, which generalizes some of the $\mathrm{Cu}$-semigroup techniques underlying Proposition 6.4 and Theorem 10.5 in [APRT24].
Proposition 5.6.
Let $(S,\Sigma)$ be a scaled $\mathrm{Cu}$-semigroup satisfying (O5) and (O7). Assume that $S$ has locally bounded comparison amplitude, and also that is upper and lower functionally almost divisible. Then $S$ is pure.
Proof.
We first prove that $S$ is almost unperforated. Let $x,y\in S$, $n\in{\mathbb{N}}$, and assume that $(n+1)x\leq ny$. We must show that $x\leq y$. This is equivalent to showing that $x^{\prime}\leq y$ for any $x^{\prime}\in S$ such that $x^{\prime}\ll x$.
Therefore, let $x^{\prime}\ll x$ and pick $x^{\prime\prime}\in S$ such that $x^{\prime}\ll x^{\prime\prime}\ll x$. Since we have $\widehat{x}\leq\tfrac{n}{n+1}\widehat{y}$, we may choose $\gamma,\gamma^{\prime}$ such that $\tfrac{n}{n+1}<\gamma<\gamma^{\prime}<1$. Using Proposition 2.3 (i) one gets $\widehat{x^{\prime\prime}}\ll\gamma\widehat{y}$, which allows us to pick $y^{\prime}\in S$ such that
$\widehat{x^{\prime\prime}}\ll\gamma\widehat{y^{\prime}},\,\,\,\text{ and }\,\,\,y^{\prime}\ll y.$
Take $y^{\prime\prime}$ such that $y^{\prime}\ll y^{\prime\prime}\ll y$, and choose $d\in{\mathbb{N}}$ such that $x^{\prime\prime},y^{\prime\prime}\in\Sigma^{(d)}$. Since $(S,\Sigma)$ has locally bounded comparison amplitude, we obtain $N=N(\gamma^{\prime},d)\in{\mathbb{N}}$ such that
$\widehat{v}\leq\gamma^{\prime}\widehat{w}\quad\Rightarrow\quad v\leq_{N}w,\quad\text{ for all }v,w\in\Sigma^{(d)}.$
Let $M\in{\mathbb{N}}$ be such that $M\geq 2$, $M\geq N$ and $\gamma\tfrac{M+1}{M-1}\leq\gamma^{\prime}$. Using (UFAD) for $x^{\prime}\ll x^{\prime\prime}$ and $M-1$, we obtain $v\in S$ satisfying
$v\leq x^{\prime\prime},\quad(M-1)\widehat{v}\leq\widehat{x^{\prime\prime}},\,\,\,\text{ and }\,\,\,x^{\prime}\leq Mv.$
Applying (LFAD) for $y^{\prime}\ll y^{\prime\prime}$ and $M$, one gets $w\in S$ such that
$Mw\leq y^{\prime\prime},\,\,\,\text{ and }\,\,\,\widehat{y^{\prime}}\leq(M+1)\widehat{w}.$
Combining the functional inequalities for $\widehat{v},\widehat{w}$ and using that $\widehat{x^{\prime\prime}}\leq\gamma\widehat{y^{\prime}}$ at the second step, we have
$(M-1)\widehat{v}\leq\widehat{x^{\prime\prime}}\leq\gamma\widehat{y^{\prime}}\leq\gamma(M+1)\widehat{w},$
which implies, by our choice of $M$, that $\widehat{v}\leq\gamma^{\prime}\widehat{w}$.
Note that both $v$ and $w$ belong to $\Sigma^{(d)}$, since $v\leq x^{\prime\prime},Mw\leq y^{\prime\prime}$, and $x^{\prime\prime},y^{\prime\prime}\in\Sigma^{(d)}$. We deduce that $v\leq_{N}w$ and, since $M\geq N$, we get $Mv\leq Mw$.
Hence,
$x^{\prime}\leq Mv\leq Mw\leq y^{\prime\prime}\leq y,$
as desired.
To show that $S$ is almost divisible, let $k\geq 2$ and take $x^{\prime},x\in S$ such that $x^{\prime}\ll x$. Since $S$ satisfies (LFAD), we may apply Lemma 4.4(ii) to obtain $y\in S$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq\left(k+\frac{1}{k}\right)\widehat{y}.$
It follows that $x^{\prime}\leq\infty y$; see, for example, [APRT22, Lemma 6.6 (iii)]. We also have that $\lambda(x^{\prime})<\lambda((k+1)y)$ for every functional normalized at $y$. Since $S$ is almost unperforated, we obtain $x^{\prime}\leq(k+1)y$ by [ERS11, Proposition 6.2]. ∎
Theorem 5.7.
Let $A$ be a $C^{*}$-algebra that has locally bounded comparison amplitude and is $n$-almost divisible for some $n\in{\mathbb{N}}$. Then $A$ is pure.
Proof.
Since $\operatorname{Cu}(A)$ is $n$-almost divisible, Theorem 4.7 gives that $\operatorname{Cu}(A)$ has (LFAD) and (UFAD), and the result follows from Proposition 5.6. ∎
Theorem 5.8.
If a $C^{*}$-algebra is $(m,n)$-pure for some $m,n\in{\mathbb{N}}$, then it is pure.
Proof.
If $A$ is $(m,n)$-pure, then since it has $m$-comparison Lemma 3.6 implies that $A$ has locally bounded comparison amplitude. Since it is also $n$-almost divisible, we may apply Theorem 5.7 to conclude that $A$ is pure. ∎
6. C*-algebras with the Global Glimm Property and finite nuclear dimension
The next result is our variation of [RT17, Proposition 3.2(ii)] for nowhere scattered $C^{*}$-algebras. This has essentially appeared in the argument of [Vil23, Proposition 4.1].
Lemma 6.1.
Let $A$ be a nowhere scattered $C^{*}$-algebra with $\dim_{\mathrm{nuc}}(A)\leq m$. Then for every $k\geq 1$, and every pair $x^{\prime},x\in\operatorname{Cu}(A)$ satisfying $x^{\prime}\ll x$, there exists $y\in\operatorname{Cu}(A)$ such that
$y\ll x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll ky\ll 2(m+1)x.$
Proof.
By [TV24, Proposition 4.12] and [WZ10, Proposition 2.3], $A\otimes\mathcal{K}$ is nowhere scattered and of nuclear dimension at most $m$. We may thus assume that $A$ is stable.
Choose $a\in A_{+}$ and $\varepsilon>0$ such that
$x^{\prime}\ll[(a-\varepsilon)_{+}],\,\,\,\text{ and }\,\,\,[a]\ll x.$
Since $A$ is nowhere scattered, the hereditary sub-$C^{*}$-algebra $\overline{aAa}$ has no finite-dimensional representations, by [TV24, Theorem 3.1]. It also has nuclear dimension at most $m$ (by [WZ10, Proposition 2.5]), whence we can apply [RT17, Proposition 3.2(ii)] for $k$ (and for $\overline{aAa}$, $a$ and $\varepsilon$) to obtain $b\in\overline{aAa}_{+}$ such that
$[(a-\varepsilon)_{+}]\leq k[b]\leq 2(m+1)[a].$
Then $y:=[b]\in\operatorname{Cu}(A)$ has the desired properties. ∎
The following is inspired by [RT17, Lemma 3.4].
Lemma 6.2.
Let $A$ be a $C^{*}$-algebra with the Global Glimm Property, and let $m,l\in{\mathbb{N}}$. Assume that $\dim_{\mathrm{nuc}}A\leq m$. Then, there exists $L=L(m,l)$ such that, for every $x^{\prime}\ll x$, there exist $y_{0},\ldots,y_{l}\in\operatorname{Cu}(A)$ such that
$y_{0}+\ldots+y_{l}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll Ly_{j}$
for $j=0,\ldots,l$.
Proof.
Let $x^{\prime},x\in\operatorname{Cu}(A)$ be such that $x^{\prime}\ll x$. We claim that there exists a separable sub-$C^{*}$-algebra $B\subseteq A$ such that
  • (i)
    $B$ has the Global Glimm Property;
  • (ii)
    $\dim_{\mathrm{nuc}}(B)\leq m$;
  • (iii)
    the induced inclusion $\operatorname{Cu}(\iota)\colon\operatorname{Cu}(B)\to\operatorname{Cu}(A)$ is an order-embedding;
  • (iv)
    $x^{\prime}$ and $x$ are in the image of $\operatorname{Cu}(\iota)$.
Indeed, we know from Proposition 2.3(iii) and 2.6 in [WZ10] that the property ‘$\dim_{\mathrm{nuc}}(\_\,)\leq m$’ is separably inheritable. Using the same methods from [TV21, Section 5], one can also see that the Global Glimm Property satisfies the same condition. Now [TV21, Proposition 6.1] allows us to choose a sub-$C^{*}$-algebra $B$ with the required conditions.
Therefore we may assume that $A$ is separable and that, as in the proof of Lemma 6.1 above, we may also assume that $A$ is stable.
Now we apply induction on $l$. Note that, if $l=0$, one can take $y_{0}:=x$. Let us assume that $l=1$.
Choose $x^{\prime\prime}\in\operatorname{Cu}(A)$ such that $x^{\prime}\ll x^{\prime\prime}\ll x$. Since $A$ has the Global Glimm Property, $\operatorname{Cu}(A)$ is $(2,\omega)$-divisible by [TV23, Theorem 3.6]. Applied for $x^{\prime\prime}$ and $x$, we obtain $c\in\operatorname{Cu}(A)$ such that
$2c\leq x,\,\,\,\text{ and }\,\,\,x^{\prime\prime}\ll\infty c.$
Choose $c^{\prime}\in\operatorname{Cu}(A)$ such that $x^{\prime\prime}\leq\infty c^{\prime}$ and $c^{\prime}\ll c$. Using Lemma 2.1 with $c^{\prime}\ll c$ and $2c\leq x$, we find $d\in\operatorname{Cu}(A)$ such that
$c^{\prime}+d\leq x\leq 2d.$
Set $e:=d\wedge\infty c^{\prime}$, which is possible by Theorem 2.2. Using at the first step that $x^{\prime\prime}\leq\infty c^{\prime}$ and $x^{\prime\prime}\leq x\leq 2d$, we get
$x^{\prime\prime}\leq(2d)\wedge\infty c^{\prime}=2(d\wedge\infty c^{\prime})=2e.$
Choose $e^{\prime\prime},e^{\prime}\in\operatorname{Cu}(A)$ such that
$x^{\prime}\ll 2e^{\prime\prime},\,\,\,\text{ and }\,\,\,e^{\prime\prime}\ll e^{\prime}\ll e.$
Applying Lemma 6.1 for $k=2m+3$ and $e^{\prime\prime}\ll e^{\prime}$, we obtain $f\in\operatorname{Cu}(A)$ such that
(2) $f\ll e^{\prime},\,\,\,\text{ and }\,\,\,e^{\prime\prime}\ll(2m+3)f\ll 2(m+1)e^{\prime}.$
Choose $f^{\prime}\in\operatorname{Cu}(A)$ such that
$e^{\prime\prime}\ll(2m+3)f^{\prime},\,\,\,\text{ and }\,\,\,f^{\prime}\ll f.$
Applying (O5) for $f^{\prime}\ll f\leq e^{\prime}$, we obtain $g\in\operatorname{Cu}(A)$ such that
(3) $f^{\prime}+g\leq e^{\prime}\leq f+g.$
Set $y_{0}:=f^{\prime}$ and $y_{1}:=g+c^{\prime}$. Then
$y_{0}+y_{1}=f^{\prime}+g+c^{\prime}\leq e^{\prime}+c^{\prime}\leq d+c^{\prime}\leq x.$
Further,
$x^{\prime}\ll 2e^{\prime\prime}\ll 2(2m+3)f^{\prime}=(4m+6)y_{0}.$
Multiplying (3) by $32m+$ 3 at the first step, and using (2) at the second step, we deduce
(4) $(2m+3)e^{\prime}\leq(2m+3)f+(2m+3)g\leq(2m+1)e^{\prime}+(2m+3)g.$
We claim that $\widehat{e^{\prime}}\leq(2m+3)\widehat{y_{1}}$. To show this, let $\lambda\in F(S)$. We distinguish two cases:
First, assume that $\lambda(c^{\prime})<\infty$. Using that $e^{\prime}\ll e\leq\infty c^{\prime}$, it follows that $\lambda(e^{\prime})<\infty$ which allows, in (4), to cancel after applying $\lambda$ to get
$\lambda(e^{\prime})\leq 2\lambda(e^{\prime})\leq(2m+3)\lambda(g)\leq(2m+3)\lambda(y_{1}).$
On the other hand, if $\lambda(c^{\prime})=\infty$, then
$\lambda(e^{\prime})\leq\infty=(2m+3)\lambda(c^{\prime})\leq(2m+3)\lambda(y_{1}).$
Since $\operatorname{Cu}(A)$ has $m$-comparison (see Remarks 3.2 (1)), we get
$e^{\prime}\leq 2m(2m+3)y_{1}.$
Using that $x^{\prime}\ll 2e^{\prime}$, one has
$x^{\prime}\ll 2e^{\prime}\leq 4m(2m+3)y_{1}.$
Thus, setting $L:=\max\{4m+6,4m(2m+3)\}$, we see that $y_{0}+y_{1}\leq x$ and $x^{\prime}\ll Ly_{0},Ly_{1}$. This shows the statement for $l=1$.
Now let $l\in{\mathbb{N}}$ and assume that we have proven the statement for $l-1$.
Take $x^{\prime},x\in\operatorname{Cu}(A)$ such that $x^{\prime}\ll x$. We get $L^{\prime}\in{\mathbb{N}}$ and elements $z_{0},\ldots,z_{l-1}$ such that
$z_{0}+\ldots+z_{l-1}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll L^{\prime}z_{j}$
for $j=0,\ldots,l-1$.
Set $y_{j}:=z_{j}$ for each $j\leq l-2$, and take $z_{l-1}^{\prime}\in\operatorname{Cu}(A)$ such that $z_{l-1}^{\prime}\ll z_{l-1}$ and $x^{\prime}\ll L^{\prime}z_{l-1}^{\prime}$. By the argument above, one finds $L^{\prime\prime}\in{\mathbb{N}}$ and $y_{l-1},y_{l}\in\operatorname{Cu}(A)$ satisfying
$y_{l-1}+y_{l}\leq z_{l-1},\,\,\,\text{ and }\,\,\,z_{l-1}^{\prime}\ll L^{\prime\prime}y_{l-1},L^{\prime\prime}y_{l}.$
One can now check that the elements $y_{0},\ldots,y_{l}$ and $L:=L^{\prime}L^{\prime\prime}$ satisfy the required conditions. ∎
Proposition 6.3.
Let $A$ be a $C^{*}$-algebra with the Global Glimm Property, and let $m\in{\mathbb{N}}$. Assume that $\dim_{\mathrm{nuc}}A\leq m$. Then, there exists $N=N(m)$ such that $\operatorname{Cu}(A)$ is $N$-almost divisible.
Proof.
The proof is inspired by that of [RT17, Lemma 3.6]. Let $L$ be the constant obtained from Lemma 6.2 for $m$ and $l=m$. We verify the statement for $N:=3L(m+1)$.
Let $x^{\prime}\ll x$ and $k\geq 1$. We need to find $y$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\leq(k+1)(N+1)y.$
Applying Lemma 6.1 for $3kL(m+1)$ and $x^{\prime}\ll x$, we obtain $c\in\operatorname{Cu}(A)$ such that
$c\ll x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll 3kL(m+1)c\ll 2(m+1)x.$
Choose $c^{\prime}\in\operatorname{Cu}(A)$ such that
$x^{\prime}\ll 3kL(m+1)c^{\prime},\,\,\,\text{ and }\,\,\,c^{\prime}\ll c.$
Now choose $x^{\prime\prime}\in\operatorname{Cu}(A)$ such that
$3kL(m+1)c^{\prime}\ll 2(m+1)x^{\prime\prime},\,\,\,\text{ and }\,\,\,x^{\prime\prime}\ll x.$
Applying Lemma 6.2 for $x^{\prime\prime}\ll x$, we obtain $y_{0},\ldots,y_{m}$ such that
$y_{0}+\ldots+y_{m}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime\prime}\leq Ly_{j}\quad\text{ for $j=0,\ldots,m$}.$
For each $j$, we obtain
$3kL(m+1)c^{\prime}\ll 2(m+1)x^{\prime\prime}\ll 2L(m+1)y_{j}.$
and thus $k\widehat{c^{\prime}}<_{s}\widehat{y_{j}}$.
By $m$-comparison, we get
$kc^{\prime}\leq y_{0}+\ldots+y_{m}\leq x.$
Further,
$x^{\prime}\ll 3kL(m+1)c^{\prime}\leq 3(k+1)L(m+1)c^{\prime}=(k+1)Nc^{\prime}\leq(k+1)(N+1)c^{\prime},$
which shows that $y=c^{\prime}$ has the desired properties. ∎
Theorem 6.4.
Every $C^{*}$-algebra with the Global Glimm Property and finite nuclear dimension is pure.
Proof.
Let $A$ be a $C^{*}$-algebra with the Global Glimm Property and that has nuclear dimension at most $m$. By Remarks 3.2 (1), we have that $A$ has $m$-comparison. Using Proposition 6.3, we know that $A$ is $n$-almost divisible for some $n$. Hence, $A$ is $(m,n)$-pure, and the result now follows from Theorem 5.8. ∎
Remark 6.5.
In [RT17, Theorem 3.1] it is shown that if a nowhere scattered $C^{*}$-algebra $A$ has finite nuclear dimension and no quotient of $A$ has a simple, purely infinite ideal, then $A$ is pure. Further, every approximately subhomogeneous $C^{*}$-algebra (ASH-algebra) has locally finite nuclear dimension [NW06]. By [ENST20, Theorem A], every $\mathcal{Z}$-stable ASH-algebra has decomposition rank at most two, and therefore finite nuclear dimension.
Combining these known results, we obtain another partial positive answer to Question B: For a separable, nowhere scattered ASH-algebra whose primitive ideal space has a basis of compact-open subsets or is Hausdorff, the properties ‘finite nuclear dimension’, ‘$\mathcal{Z}$-stability’ and ‘pureness’ are equivalent. If one additionally assumes that the algebra has stable rank one, then these properties are also equivalent to ‘strict comparison of positive elements’.